# MALADY: Multiclass Active Learning with Auction Dynamics on Graphs

### v2

@@ 3.3 Multiclass Active Learning Procedure
We now describe our proposed Multiclass Active Learning with Auction DYnamics on graphs (MALADY) procedure. Given an initially labeled set $\mathcal{L}_{0}$ with corresponding labels $\{y(x_{i})\}_{x_{i}\in\mathcal{L}}$, we sequentially select query points that maximize our proposed acquisition function (26), $\mathcal{A}(x)=1-\mathcal{M}(x)$. We summarize our method as follows: Given input data and an initially labeled subset $\mathcal{L}_{0}$, our proposed method consists of the following steps:
@@
    Using the acquisition function (26), select the (next) query point as its maximizer: $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$.
@@ 4.1 Data sets
For the computational experiments in this paper, we have used six benchmark data sets, including two hyperspectral data sets. The data sets are the following:
  • The Landsat data set [47] is a data set of 6435 elements containing multi-spectral values of pixels in 3x3 neighborhoods in a satellite image. For this data set, the classification is associated with the central pixel and there are six classes. Each of the data elements contains the pixel values in the four spectral bands of each pixel in the 3×3 neighborhood. Therefore, each feature vector consists of 36 values. The six classes are red soil, cotton crop, grey soil, damp grey soil, soil with vegetation stubble, and very damp grey soil.
  • The USPS data set [48] is a data set of 9,298 square grayscale images that contain 16×16 pixels, using a 4-bit grey scale system with 256 gray levels per pixel. The images were obtained by automatically scanning envelopes by the United States Postal Service in Buffalo, New York. The grayscale images are centered and normalized and show various font styles. There are 10 classes; each data element is an image of a handwritten digit “0” through “9”.
  • The Jasper Ridge data set [49] is one of the most widely used hyperspectral data sets in the hyperspectral image classification task. It consists of an $512\times 614$ image with 198 channels. For our experiments, we only consider a subimage of the image consisting of $100\times 100$ pixels with 198 channels.
  • The Urban data set [50] is one of the most widely used hyperspectral data sets in hyperspectral applications. It consists of an $307\times 307$ image with 162 channels. The version of the data set we use constains four endmembers corresponding to asphalt, grass, tree and roof. The spectral resolution is 10 nm, and each pixel corresponds to an area of two square meters. The data set was recorded in 1995 by the Hyperspectral Digital Imagery Collection Experiment in Copperas Cove, Texas.
  • The COIL-20 data set [51] is a multiclass data set of 1440 normalized images of 20 objects. The objects were placed on a motornized turn table. With a fixed camera, the turn table was rotated through 360 degrees to capture different poses every 5 degrees. There are a total of 72 images per object.
  • The Opt-Digits data set [52] is a multiclass data set of 56200 handwritten digits. Each handwritten digit is recorded as a 8 $\times$ 8 matrix. Thus, each element has 64 attributes.
The details of the data sets are outlined in Table 1.
Table 1: Data sets used in the experiments.
Data set Number of elements Number of attributes Number of classes
Coil-20 [51] 1400 1024 20
Opt-Digits [52] 5620 64 10
Landsat [47] 6435 36 6
USPS [48] 9298 64 10
Jasper Ridge [49] 10000 198 4
Urban [50] 94249 162 4
@@
Figure 4: Accuracy v.s. # of labeled points for each dataset: (a) Optdigits (b) USPS (c) Satellite (d) COIL-20 (e) Urban (f) Jasper Ridge
@@ 4.2 Comparison to other methods
In this section, we compare our proposed technique to those using other active learning acquisition functions in various graph-based semi-supervised methods. The results of the experiments and the comparison methods are shown in Figure 2. In particular, for all data sets, we compare our proposed method, MALADY, to the following methods.
  • Random sampling (Random)
  • Variance minimization criterion (VOpt) [5]
  • Minimum norm uncertainty sampling (Unc. (Norm))[4]
  • Uncertainty sampling (Unc-Laplace) [1]
  • k-means++ [53]
For a fair comparison, we used Algorithm 1 as the underlying semi-supervised method for the Random, V-OPT and k-means++ algorithms. For minimum norm uncertainty sampling, we used Poisson reweighted Laplace learning [54] as the underlying method. For Unc-Laplace, the underlying classifier is Laplace learning [55].
@@
Now, we provide a brief overview of the methods that we include in our numerical comparisons. In the V-Opt algorithm [5], the authors analyze the probability distribution of the unlabeled vertices conditioned on the label information, which follows a multivariate normal distribution with the mean corresponding to the harmonic solution across the field. The nodes are then selected for querying such that the total variance of the distribution on the unlabeled data, as well as the expected prediction, is minimized. In minimum norm uncertainty sampling [4], the authors design an acquisition function that measures uncertainty in the Poisson reweighed Laplace learning algorithm (PWLL). In particular, the authors control the exploration versus exploitation tradeoff in the active learning process by introducing a diagonal perturbation in PWLL which produces exponential localization of solutions. The uncertainty sampling-based method [1] queries points near decision boundaries that are most uncertain. Common uncertainty measures include least-confident, smallest margin, and entropy. For comparison, we use the smallest margin acquisition function with Laplace learning as an underlying classifier, which prioritizes points near the decision boundaries. Lastly, we use the seeding method for k-means++ [56] which chooses an initial set of diverse centroids for the k-means clustering algorithm [57] to select centroids within each distinct cluster. One can think of labeling these initial centroid sets to obtain representative and diverse labeled sets within an active learning framework, and we therefore include it as a method for comparison.
@@
For the implementation of MALADY, V-OPT, Random, k-means++, and Unc-Laplace, we utilized the GraphLearning Python package [19] which contains both graph learning and active learning frameworks. For the implementation of Unc. (Norm) [4], we utilize the code obtained from the author’s repository.
@@ 4.3 Hyperparameters Selection
In this section, we outline the selected hyperparameters for our experiments. In particular, the number of nearest neighbors, $N_{n}$, for the weight matrix is a tunable parameter that can be adjusted to control the sparsity of the graph. For the computational experiments, we set the number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, and the initial epsilon $\epsilon_{0}$ to 100, $1e^{-6}$, 4, and $1e^{-7}$, respectively. We also incorporate class size constraints for the classification which are controlled by the upper bound $U$ and lower bound $B$. In our computational experiments, we use exact class size constraints. Overall, we have outlined all parameters used for the computational experiments in the Supplementary Information file.
@@ 4.4 Performance and Discussion
Information about the data sets used for our computational experiments is detailed in Section 4.1 and Table 1. For each data set, we use five initial labels per class; therefore, the initial labeled set has a total of $5N$ labeled points, where $N$ is the number of classes. Then, we sequentially query 500 additional data elements to label. Moreover, for both hyperspectral datasets, Urban and Jasper Ridge, we use the cosine similarity function (2) for graph weight computation, and we use the Gaussian weight function (1) for the other data sets.
In Figure 4, we show the accuracy performance of each acquisition function averaged over 10 trials, except for the Urban dataset, where we averaged over 5 trials; for all data sets, accuracy was the main evaluation metric. From the results, one can see that our MALADY algorithm has significantly outperformed the other comparison methods in all data sets except COIL-20. For the COIL-20 data set, our MALADY algorithm performs similarly to the best-performing method Unc. Norm [4] which uses a different underlying graph-based classifier than our auction dynamics model.
@@ 5 Conclusion
In this paper, we introduce the Multiclass Active Learning with Auction DYnamics on Graphs (MALADY) algorithm which integrates active learning with auction dynamics techniques for semi-supervised learning and data classification and ensures exploitation in the active learning process. The proposed method allows one to obtain accurate results even in the case of very small labeled sets, a common scenario for many applications since it can be costly in time and money to obtain labeled data in many applications. The proposed method leverages a novel acquisition function derived from the dual variables in an auction algorithm. Lastly, the proposed algorithm also allows one to incorporate class size constraints for the data classification task, which improves accuracy even further. Overall, the proposed procedure is a powerful approach for an important machine learning task.

### v3

@@ 3.3 Multiclass Active Learning Procedure
We now describe our proposed Multiclass Active Learning with Auction DYnamics on graphs (MALADY) procedure. Given an initially labeled set $\mathcal{L}_{0}$ with corresponding labels $\{y(x_{i})\}_{x_{i}\in\mathcal{L}}$, we sequentially select query points that maximize our proposed acquisition function (26), $\mathcal{A}(x)=1-\mathcal{M}(x)$. We summarize our method as follows: Given input data and an initially labeled subset $\mathcal{L}_{0}$, our proposed method consists of:
@@
    Using acquisition function (26), select the (next) query point as: $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$.
@@ 4.1 Data sets
For the computational experiments in this paper, we use eight benchmark datasets, including two hyperspectral datasets: Landsat [57], USPS [58], Jasper Ridge [59], Urban [60], COIL-20 [61], CIFAR-10 [62], Opt-Digits [63], and Fashion-MNIST [64]. The details of these data sets are provided in the supporting documentation.
@@
g) Refer to caption
h) Refer to caption
Figure 4: Accuracy v.s. # of labeled points for each dataset: (a) Opt-Digits (b) USPS (c) Satellite (d) Jasper Ridge (e) Urban (f) COIL-20 (g) CIFAR-10 (h) FASHIONMNIST. To allow for a more fair comparison of each acquisition function’s selections for labeled sets, we report accuracy results for Random, VOpt, MCVOpt, and $S^{2}$ using Algorithm 1 since these acquisition functions are not tied to a particular graph-based semi-supervised classifier. All other methods display accuracy results from their corresponding graph-based SSL classifiers.
@@ 4.2 Comparison to other methods
In this section, we compare our proposed technique to those using other active learning acquisition functions in various graph-based SSL classifiers. The results of the experiments are shown in Figure 2. In particular, for all data sets, we compare our proposed method, MALADY, to the following methods: Random sampling (Random), Variance minimization criterion (VOpt) [5], Minimum norm uncertainty sampling (Unc.-Norm)[4], Model change active learning (MCVOpt) [65], Uncertainty sampling (Unc.-Laplace) [1], and $\text{S}^{2}$ [66].
For a fair comparison, we used Algorithm 1 as the underlying SSL classifier for the Random, VOpt, MCVOpt, and $\text{S}^{2}$ methods since these acquisition functions are not tied to a particular, underlying graph-based semi-supervised classifier. Minimum norm uncertainty sampling (Unc.-Norm) and Unc.-Laplace, however, are acquisition functions that at each iteration depend upon the underlying classifier’s predictions to quantify uncertainty, so we report the accuracy in their originally proposed graph-based SSL classifiers; respectively, Poisson-reweighted Laplace learning [67, 4] and Laplace learning [55].
Now, we provide a brief overview of our comparison methods. In [5] (VOpt method), the authors analyze the probability distribution of unlabeled vertices conditioned on the label information, which follows a multivariate normal distribution with mean corresponding to the harmonic solution across the field. The nodes are selected for querying such that the total variance of the distribution on the unlabeled data, as well as the expected prediction, is minimized. The MCVOpt (Model Change plus VOpt) acquisition function [68] is a heuristic combination of Model Change [6, 7] and VOpt [5] acquisition functions. The Model Change [6] acquisition function computes the change in a soft-constraint Laplace learning classifier that a currently unlabeled point could induce as a result of labeling it; this is approximated with a spectral truncation (low-rank approximation) of the graph Laplacian to improve efficiency.
@@
In minimum norm uncertainty sampling [4], the authors design an acquisition function that measures uncertainty in the Poisson reweighed Laplace learning algorithm (PWLL). In particular, the authors control the exploration versus exploitation tradeoff in the active learning process by introducing a diagonal perturbation in PWLL which produces exponential localization of solutions. The uncertainty sampling-based method [1] queries points near decision boundaries that are most uncertain. Common uncertainty measures include least confidence, smallest margin, and entropy. For comparison, we use the smallest margin acquisition function with Laplace learning as an underlying classifier, which prioritizes points near the decision boundaries. It should be noted that our proposed acquisition function for MALADY can be thought of as a form of uncertainty sampling where the quantification of uncertainty is derived from the auction dynamics graph-based classifier (Algorithm 1). Finally, the $\text{S}^{2}$ [66] algorithm queries the label of the vertex that bisects the shortest-shortest path between any pair of oppositely labeled vertices.
@@
For the implementation of MALADY, VOpt, MCVOpt, Random, and Unc.-Laplace, we utilized the GraphLearning Python package [19] which contains both graph learning and active learning frameworks. For the implementation of Unc.-Norm [4] and $S^{2}$, we utilize the code obtained from the author’s repository.
@@ 4.3 Hyperparameters Selection
In this section, we outline the selected hyperparameters for our experiments. In particular, the number of nearest neighbors, $N_{n}$, for the weight matrix is a tunable parameter that can be adjusted to control the sparsity of the graph. For the computational experiments, we set the number of steps $s=100$, auction error tolerance $\epsilon_{min}=1e^{-6}$, epsilon scaling factor $\alpha=4$, and the initial epsilon $\epsilon_{0}=1e^{-7}$. We also incorporate class size constraints for the classification, which are controlled by the upper bound $U$ and lower bound $B$. In our computational experiments, we use exact class size constraints. See the Supplementary Information file for a comprehensive outline of all parameters used in the experiments, along with a discussion of their respective roles.
4.4 Ablation Studies
In order to evaluate the effectiveness and robustness of our method, we perform ablation studies on the role of the class size constraint, the initial labeled set, and the proposed acquisition function. Firstly, to examine the performance of our method with various class size constraints, we run the experiments in two datasets using various bounds. In particular, we test USPS and Opt-Digits with various class size constraints and show the result in Figure (5). We observe that our method performs better with exact constraints, especially in a low labeled rate. Second, to assess the effectiveness of the auction dynamics algorithm, we ablate on the semi-supervised learning (SSL) classifier for output prediction accuracy on unlabeled sets. Specifically, we use our MALADY framework to sample the labeled set and tested different SSL classifiers to evaluate model accuracy. The results, presented in Figure 6, compare the performance of Multiclass MBO [30], Poisson learning [19], Laplace learning [55], and Poisson-reweighted Laplace learning (PWLL) [67]. The results in Figure 6 show that our proposed acquisition function in MALADY achieves optimal performance when paired with the auction dynamics algorithm (Algorithm 1) as the underlying SSL classifier. We also conduct additional ablation studies to understand how our framework performs with varying numbers of initially labeled samples and different distributions. Due to page limitations, however, we provide this ablation study’s results in the Supporting Information file.
a) Refer to caption
b) Refer to caption
Figure 5: MALADY under various class size constraints. The red curve represents the exact class size constraint, the black curve represents the flexible constraint, and the pink curve represents no class size constraint. (a): USPS data set (b): Opt-Digits data set.
a) Refer to caption
b) Refer to caption
Figure 6: Performance of different SSL classifiers on labeled set sampled using our MALADY framework. (a): USPS data set (b): Landsat data set.
@@ 4.5 Performance and Discussion
Information about the data sets used for our computational experiments is detailed in Supporting Information file. For each data set, we use five initial labels per class; thus, the initial labeled set has a total of $5K$ labeled points, where $K$ is the number of classes. We sequentially query 500 additional data elements to label. Moreover, for the hyperspectral data, we use the cosine similarity function (2) for graph weight computation. For others, we use the Gaussian weight function (1).
In Figure 4, we show the accuracy performance of each acquisition function averaged over 10 trials; for all data sets, accuracy was the main evaluation metric. From the results, one can see that our MALADY method has significantly outperformed the comparison methods in all data sets except COIL-20. For the COIL-20 data set, our MALADY algorithm performs similarly to the best-performing method Unc.-Norm [4], which uses a different underlying graph-based classifier than our model.
The significant difference in performance of our proposed MALADY method for active learning over other uncertainty-based acquisition functions (e.g., Unc.-Norm and Unc.-Laplace) warrants some further discussion. The success of uncertainty sampling active learning methods most often depends on the capability of the underlying classifier to use small amounts of labeled data to produce reasonable decision boundaries that can be highlighted by the corresponding acquisition function. Recall that our proposed acquisition function utilizes the auction values (i.e., coefficients, prices, and incentives) computed by the auction dynamics classifier (Algorithm 1) to identify the boundaries between clusters containing oppositely labeled nodes. Both the extensive empirical results (Figure 4) and the intuitive toy example visualized (Figure 2) suggest that these auction values are quite useful for accurately modeling the geometry of the clusters, especially these corresponding decision boundaries even with low amounts of labeled data. While a rigorous theoretical analysis of the mechanism of Algorithm 1 that influences such accurate modeling of clustering geometry is outside the scope of the current work, we humbly suggest that this graph-based auction dynamics procedure is a key distinguishing component of MALADY that allows for superior uncertainty sampling performance.
@@ 5 Conclusion
In this paper, we introduce the Multiclass Active Learning with Auction DYnamics on Graphs (MALADY) method which integrates active learning with auction dynamics techniques for semi-supervised learning and classification and ensures exploitation in the active learning process. The proposed method allows one to obtain accurate results even in the case of very small labeled sets, a common scenario for many applications since it can be costly in time and money to obtain labeled data in many applications. The proposed method leverages a novel acquisition function derived from the dual variables in an auction algorithm. Lastly, the proposed algorithm also allows one to incorporate class size constraints for the classification task, which improves accuracy even further. Overall, the proposed procedure is a powerful approach for an important machine learning task.
Acknowledgements
This work is supported by NSF grant DMS-2052983.